### v2

# Dynamical Lensing Tomography of Black Hole Ringdown

Abstract
Strong gravitational lensing occurs when photons pass through the vicinity of a black hole. We investigate this phenomenon in the context of a gravitational-wave event, specifically when a black hole is settling into its final state. The deflection angle of photons mimics the ringdown pattern of the gravitational wave at intermediate times. At late times it has an inverse cubic dependence on observation time. The deviation angle increases exponentially as photons approach the photon ring orbit, reflecting its unstable nature. Our findings are directly applicable to imaging scenarios involving stars against the background of compact binaries, or circumbinary accretion disks, particularly during the merger of two black holes.
I Introduction
One of the foundational tests of General Relativity concerns the deflection of light by massive bodies Will (2014); Lemos et al. (2019); Bartelmann and Schneider (2001); Bartelmann (2010): the path of light rays is “bent” by mass-energy, which distorts the spacetime fabric. In the last decade, it has become clear that the universe is flooded with traveling spacetime distortions, gravitational waves (GWs) traveling at the speed of light. Gravitational-wave astronomy provides a wealth of information on compact objects, particularly on the behavior of gravity in highly dynamical, high-curvature and redshift setups.
An intriguing prospect concerns the detectability of light bending by GWs, which would add an extra information channel on strong gravity. This possibility was raised and studied by different authors in the weak field regime Labeyrie (1993); Fakir (1994); Durrer (1994); Kaiser and Jaffe (1997); Damour and Esposito-Farese (1998). Our purpose here is to study light deflection in the strong-field, highly dynamical regime. In particular, we will study the appearance of a black hole (BH) which is approaching a quiescent state via relaxation in its characteristic modes of oscillation, called quasinormal modes (QNMs). The relaxation stage is also referred to as ringdown stage, and is a generic late-time description of the coalescence of compact objects Berti et al. (2009); Baibhav et al. (2023). Our results could describe stars on the background of a binary BH merger, or the appearance of a circumbinary accretion disk surrounding two merging compact objects.
We work in units where $G=c=1$, and adopt the metric convention (,+,+,+(-,+,+,+)( - , + , + , + ).
II Ray Tracing towards a Relaxing Black Hole
A relaxing BH is well described by general relativistic first-order perturbation theory, where spacetime is expressed by small fluctuations away from the final, stationary state. Uniqueness results in General Relativity suggest that the final state, to a good approximation, is a Kerr BH Chrusciel et al. (2012); Cardoso and Gualtieri (2016); Cardoso et al. (2022). We thus take our relaxing BH spacetime, of mass $M$ and angular momentum $aM$ to be given by
$g_{\mu\nu}=g_{\mu\nu}^{\mathrm{(0)}}+\,h_{\mu\nu}\,,$ (1)
where $g_{\mu\nu}^{\mathrm{(0)}}$ is the Kerr metric (we will use Boyer-Lindquist $(t,r,\theta,\phi)$ coordinates throughout Bardeen et al. (1972)). The term $h_{\mu\nu}$ denotes first-order perturbations.
Instead of working directly with the metric perturbations $h_{\mu\nu}$, we consider instead curvature perturbations Teukolsky (1973, 1972). These can be separated and decoupled and all the relevant information about radiative degrees of freedom is encapsulated in a master variable
$\psi_{s}(t,r,\theta,\phi)=\frac{{}_{s}Z}{\sqrt{2\pi}}\,{}_{s}S(\theta)\,{}_{s}R(r)e^{-i\omega t+im\phi}\,.$ (2)
The gravitational quantity of interest is $\Psi_{4}^{(1)}=\psi_{-2}(t,r,\theta,\phi)(r-ia\cos\theta)^{-4}$, the first-order perturbation of one of the Weyl scalars, which is directly related to the amplitude of GWs far from the source. Notation and conventions are shown in Supplemental Material A 1. The function ${}_{s}S(\theta)$ is a spin-weighted spheroidal harmonic, while the radial Teukolsky function ${}_{s}R(r)$ obeys a homogeneous second-order differential equation. We fix its normalization by insisting on the large-distance behavior. Then, ${}_{s}Z$ controls the amplitude of the GW, by matching the ringdown stage to some physically generated process. The characteristic modes are found by imposing proper boundary conditions on ${}_{s}R(r)$ , which selects only a discrete set of frequencies $\omega$, the ringdown frequencies Berti et al. (2009). All the functions and quantities carry a dependence on angular numbers $\ell,m$ and a possible overtone number $n$ Teukolsky (1973, 1972); Berti et al. (2009). For simplicity, we do not include these indices in the subscripts. Although a true GW signal comprises a superposition of various modes, our discussion focuses on the dominant $\ell=m=2$ mode. Additionally, we exclusively address the $n=0$ mode, which lasts longer compared to the more rapidly decaying higher overtones. For QNMs characterized by frequency $\omega$ and labeled by $(\ell,m,n)$, there exists a corresponding “mirror” mode with frequency $-\bar{\omega}$ and $(\ell,-m,n)$. In the final spin frame, these paired modes primarily represent GW emissions propagating in opposite directions along the spin axis—one mode predominantly radiating towards the northern hemisphere and the other towards the southern hemisphere Li et al. (2022); Berti et al. (2006); Cook and Zalutskiy (2014); Krivan et al. (1997). The observed waveform typically manifests as a superposition of these two distinct sets of modes. However, one of the paired modes may be less prominent due to weaker excitation Berti et al. (2006); Hamilton et al. (2023). In the following analysis, we primarily focus on the $m=2$ mode with $\mathrm{Re}(\omega)>0$, which predominantly emits in the southern hemisphere, as more general cases exhibit similar features.
Once ${}_{-2}R(r)$ is known, we reconstruct the metric fluctuation $h_{\mu\nu}$ following well-known prescriptions Chrzanowski (1975); Cohen and Kegeles (1974); Kegeles and Cohen (1979), summarized in Supplemental Material A 2. The QNMs diverge as $h_{\mu\nu}\sim e^{\mathrm{Im}(\omega)(t-r_{*})}$ when $t-r_{*}\to-\infty$, where $\mathrm{Im}(\omega)$ represents the imaginary part of the ringdown frequency $\omega$ and $r_{*}$ is the tortoise coordinate Berti et al. (2009). To mitigate the divergence, we truncate the perturbation metric $h_{\mu\nu}$ by applying a Heaviside-like filter function $H(t-r_{*})$:
$H(t-r_{*})=\frac{1}{2}\left(\tanh{\frac{t-r_{*}}{q}}+1\right)\,,$ (3)
to obtain $H(t-r_{*})\,h_{\mu\nu}$. Here, $0<q<-2/\mathrm{Im}(\omega)$, where $q$ controls the intermediate region width of the filter. Although various values of $q$ were tested, the modifications had only a minor impact on the outcome, especially regarding the late-time frequency behavior. In the following, we employ a relatively sharp truncation in the filter function $H(t-r_{*})$, with $-q\mathrm{Im}(\omega)=0.1$.
During the ringdown phase, a photon trajectory is influenced by the metric perturbations encountered along its path, governed by the geodesic equation:
[…]
where $\Gamma$ denotes the Levi-Civita connection, and $\tau$ represents the proper time.
To characterize the observable effects, it is essential to define the observer frame. In an unperturbed Kerr BH scenario, a zero angular momentum observer (ZAMO) frame is typically used Frolov and Novikov (1998). However, in a dynamic spacetime, the ZAMO tetrads are neither orthogonal nor normalized. To rectify this, we apply a Gram-Schmidt orthogonalization process to establish a normalized, orthogonal tetrad basis derived from the ZAMO tetrads, as detailed in Supplemental Material B 1. Using this newly defined observer frame, we initiate geodesics towards the BH employing backward ray tracing. Here, each pixel on the observer plane, represented by coordinates $(\tilde{X},\tilde{Y})$ Bardeen (1973); Gralla et al. (2018); Gralla and Lupsasca (2020), corresponds to photon momentum in this frame. The photon’s 44 44-momentum in the dynamic spacetime $p^{\mu}\equiv{\rm d}x^{\mu}/{\rm d}\tau$ is then constructed from this orthogonal tetrad basis, similar to the procedure established in Kerr spacetime Cunha et al. (2016). For convenience, we also introduce polar coordinates for the observer plane centered on the direction directly facing the BH in Kerr spacetime, denoted as $(\tilde{\rho},\tilde{\varphi})$, where $\tilde{\rho}$ is the impact parameter and $\tilde{\varphi}$ is the angle.
Photon geodesics in the relaxing Kerr spacetime are then computed. We position an observer at $r_{o}=10^{9}\,M$, confirming convergence of results at large $r_{o}$ due to the decay of $h_{\mu\nu}$ amplitude as $1/r$ at greater distances. Notice that even though we employ backward ray tracing, photon propagation in the $+t$ direction still aims towards the observer, aligning with the GWs that co-propagate in the same direction upon departure from the BH.
III Lensing Tomography
Figure 1: Top: Illustration of photon trajectories in a relaxing Kerr BH spacetime with spin $a=0.7M$, viewed from a face-on observer with $\theta_{0}\approx 0$. The metric perturbation is modeled using the $n=0$ and $\ell=m=2$ mode, with magnitude normalized to $|{}_{-2}Z|=0.24$. The initial impact parameters are set at $\tilde{\rho}=5.041M$, slightly exceeding the critical curve $\tilde{\rho}_{c}=5.04M$ in Kerr spacetime. The black line represents unperturbed geodesics in Kerr spacetime, an arrow indicating the direction from the observer to the BH to depict the backward ray tracing method used. Different colors represent geodesics launched at various observation times $t_{0}-r_{*}(r_{0})$, detailed in the bottom panel. The dashed line marks a trajectory terminating on the BH, whereas the others extend towards infinity. The gray plane indicates the BH equatorial plane. Bottom: Growth of geodesic deviation $\Delta_{\rm KS}$ representing the distance between unperturbed and deflected geodesics at the same backward propagating time $t_{0}-t-r_{*}(r_{0})$, shown in Cartesian KS coordinates.
We explore geodesics in a relaxing Kerr spacetime, particularly focusing on the merger events of supermassive binary BHs (SMBHBs) with equal mass ratios. The remnant BH from each mass ratio merger events typically characterized by an angular momentum parameter $a\sim 0.7\,M$ Buonanno et al. (2007), a value consistent with numerous observations Nitz et al. (2019); Abbott et al. (2019) and considered as our fiducial value. The GW signal from these mergers is dominated by the $\ell=|m|=2$ mode Berti et al. (2007, 2009, 2006); Maggiore (2018), with the $n=0$ mode persisting the longest. This fundamental mode is characterized by $M\omega=0.5326-0.08079i$ Berti et al. (2009); Stein (2019).
In the vicinity of the BH, the metric perturbation amplitude can reach up to $M^{2}\Psi_{4}^{(1)}\sim 0.1$ Berti et al. (2007). The total radiated energy during the ringdown is limited to approximately $3\%$ of the remnant BH’s mass Berti et al. (2007), setting an upper limit on the magnitude of $|{}_{-2}Z|$ for the $\ell=m=2$ mode at about 0.240.24 0.240.24. We adopt this upper bound as the fiducial value for our analyses, our results can be trivially re-scaled for generic amplitudes.
We first consider a face-on observer with $\theta_{o}\approx 0$. The top panel of Fig. 1 illustrates backward photon trajectories, all launched from the same direction, using identical backward impact parameters relative to the observer positioned at $r_{0}=10^{9}M$. These trajectories are depicted using the Cartesian Kerr–Schild (KS) coordinate system $(x,y,z)$. The black line represents geodesics in an unperturbed Kerr spacetime, while various colors indicate different emission times $t_{o}-r_{*}(r_{o})$. The bottom panel displays the growth of deviations between unperturbed and deflected geodesics, defined as $\Delta_{\rm KS}\equiv(\Delta x^{2}+\Delta y^{2}+\Delta z^{2})^{1/2}$, representing the distance between the two trajectories at the same backward propagating time $t_{0}-t-r_{*}(r_{o})$, colors corresponding to trajectory lines in the top panel.
In Fig. 1, we select $\tilde{\rho}=5.041M$ as the (backward) impact parameter, which is close to the critical curve $\tilde{\rho}_{c}\approx 5.04M$ for this BH spin. This curve, defined in Kerr spacetime, represents the geodesics that can propagate around the BH indefinitely. Some of the perturbed geodesics, shown in dashed lines, terminate on the BH instead of propagating towards infinity, indicating a distortion of the critical curve during the ringdown.
The evolution of geodesic deviations, launched at different observer times, initially shows a weak increase after departing from the observer. Geodesics launched later exhibit deviations exponentially smaller than those launched earlier, reflecting the exponential decay of GW amplitude near the BH during the ringdown phase. Closer to the BH, all deviations begin to exhibit exponential growth due to the instability of the radial potential in the light ring region Podurets (1965); Ames and Thorne (1968); Luminet (1979); Campbell and Matzner (1973); Cunningham and Bardeen (1972); Falcke et al. (2000); Chandrasekhar (1983); Cardoso et al. (2009); Johnson et al. (2020); Gralla and Lupsasca (2020); Cardoso et al. (2021); Chen et al. (2023). Here, the interplay between light ring instability and GW waves induces addiontial oscillation features in the exponential growth. The geodesics that depart from the light ring region begin to develop deviations that increase linearly with time.
To quantify the lensing during ringdown, we analyze backward photons that ultimately reach infinity, corresponding to far-away sources. We assess the angular difference between the asymptotic direction for Kerr geodesic and the deflected geodesic as observables, including the grand arc difference $\tilde{\delta}$, and variations in the azimuthal angle $\Delta\phi$ and polar angle $\Delta\theta$.
Figure 2: Variations of the deflected angles in the azimuthal ($\Delta\phi$) and polar ($\Delta\theta$) components for geodesics terminating at infinity, plotted as a function of observation time $t_{0}-r_{*}(r_{o})$. The BH, observer, and ringdown configuration follow the setup described in Fig. 1. Different colors represent geodesics launched from various angles $\tilde{\varphi}$ on the polar coordinate of the observer plane, with impact parameters fixed at $\tilde{\rho}=5.4M$. Variations for cases separated by $\pi$ in $\tilde{\varphi}$ appear identical, reflecting the $m=2$ nature of the BH ringdown.
Figure 2 displays the deflected angles in polar $(\Delta\theta)$ and azimuthal $(\Delta\phi)$ components as a function of observer frame time, $t_{0}-r_{*}(r_{o})$, for trajectories initially launched from various polar coordinate angles $\tilde{\varphi}$ on the observer plane. The impact parameters are fixed at $\tilde{\rho}=5.4M$. Due to ringdown metric perturbations, axial symmetry is not preserved even in face-on observations. Notably, both $\Delta\phi$ and $\Delta\theta$ show identical variations for values of $\tilde{\varphi}$ separated by $\pi$. This pattern is precisely attributable to the $m=2$ mode of the metric perturbations, akin to those generated by vector or tensor superradiant clouds Chen et al. (2023), demonstrating that the deflected angles can effectively dissect the polarization pattern of the GWs around the BH.
Next, we adjust the impact parameters $\Delta\tilde{\rho}\equiv\tilde{\rho}-\tilde{\rho}_{c}$ relative to the critical curve $\tilde{\rho}_{c}\approx 5.04M$, as defined in Kerr spacetime. In Fig. 3, we illustrate variations of the deflected angle $\tilde{\delta}$. Due to the implementation of the filter, deflection begins around $t_{0}-r_{*}(r_{o})\approx 0$. For impact parameters close to the critical curve, specifically in cases where $\Delta\tilde{\rho}\leq 1.5M$, the early time evolution of the deflected angle exhibits a pattern of exponential decay and oscillation. The timescales of this decay and the oscillation period align with the inverses of the imaginary and real parts of $\omega$, respectively, suggesting that these observables can effectively capture key characteristics of the BH ringdown.
Furthermore, when comparing the overall amplitude of $\tilde{\delta}$, there is a noticeable exponential increase dependent on $\Delta\tilde{\rho}$. This enhancement is again attributed to the instability of the BH light ring region, which causes an exponential growth in the perturbed deviation during propagation around the BH Podurets (1965); Ames and Thorne (1968); Luminet (1979); Campbell and Matzner (1973); Cunningham and Bardeen (1972); Falcke et al. (2000); Chandrasekhar (1983); Cardoso et al. (2009); Johnson et al. (2020); Gralla and Lupsasca (2020); Chen et al. (2023). As $\tilde{\rho}$ approaches the critical curve, this instability intensifies. The instability is quantified by the square root of the second derivative of the radial potential Johnson et al. (2020), which increases linearly as $\Delta\tilde{\rho}$ decreases, thereby leading to an exponential rise in the deflected angle.
Figure 3: Deflection angle of geodesics terminating at infinity, defined as the grand arc difference with the unperturbed direction, plotted as a function of the observation time $t_{0}-r_{*}(r_{o})$. Different colors represent geodesics launched with various impact parameter shifts $\Delta\tilde{\rho}\equiv\tilde{\rho}-\tilde{\rho}_{c}$ from the critical curve in Kerr spacetime, where $\tilde{\rho}_{c}\approx 5.04M$, along the $+\tilde{X}$-axis. The BH, observer, and ringdown configuration are consistent with the setup described in Fig. 1. Near the critical curve, we find an oscillatory, exponentially decaying dependence at intermediate times, in very good (to sub-percent level) agreement with the ringdown of the GW itself. At later times, a pattern of slow decay is observed, consistent with a power law $\tilde{\delta}\sim(t_{o}-r_{*})^{-p}$ with $p\sim 3$.
On the other hand, for late behavior and cases with large impact parameters $\Delta\tilde{\rho}/M=5,10,15$, the variations exhibit a slow decay without oscillations. The two different types of behavior is attributable to the deflections developing in two distinct regions. From the BH to the observer in the $+t$ direction, the GW and photons co-propagate, so the deflection that develops along this path reflects the ringdown metric perturbations present when the photon starts to co-propagate with the GWs. Before entering the co-propagation region, particularly as the photon travels from infinity towards the BH, the deflection received along the path is predominantly influenced by the time it passes through the filter at $t\approx r_{*}$. For smaller impact parameters, the co-propagation duration lasts longer, resulting in deflected angles starting with an exponential decay and oscillation, mirroring the GW ringdown. However, in late-time scenarios and with larger impact parameters, the deflection is primarily influenced by non-co-propagating impacts after $t\approx r_{*}$. The case with $\Delta\tilde{\rho}=15M$ shows larger deflection than those with $\Delta\tilde{\rho}=5M$ and $10M$ because its polar angle $\theta$ at $t\approx r_{*}$ is closer to $\pi$, where the metric perturbation for the $\ell=m=2$ mode is most dominant. Our results are consistent with a power law, $\tilde{\delta}\sim(t_{o}-r_{*})^{-p}$, where $p\approx 3$. This scaling behavior resembles that observed in GW lensing from localized sources Damour and Esposito-Farese (1998); Kopeikin et al. (1999); Kopeikin and Schaefer (1999); Kopeikin et al. (2006), where the dominant contribution to lensing at distances far from the source is proportional to $h_{\mu\nu}p^{\mu}p^{\nu}$, which is only non-vanishing for the $\mathcal{O}(1/r^{3})$ component of the metric perturbations.
Figure 4: Similar to Fig. 3, this figure displays variations in the deflected angles of geodesics terminating at infinity, as observed from different inclination angles $\theta_{0}$ and plotted against observation time $t_{0}-r_{*}(r_{o})$. The impact parameter is fixed at $\tilde{\rho}=10M$ on the $+\tilde{X}$ axis.
We also display cases with different observer inclination angles $\theta_{o}$ in Fig. 4, fixing the impact parameters at $\tilde{\rho}=10M$ along the $+\tilde{X}$-axis. Similar to the observations in Fig. 3, the variations exhibit two types of behavior: a slow decay predominates at smaller $\theta_{o}$, while larger $\theta_{o}$ are primarily characterized by ringdown-like decay and oscillation. This distinction can again be attributed to the propagation of the polar angle from infinity to the BH, where smaller $\theta_{o}$ corresponds to larger $\theta$ in this trajectory. For $\theta_{o}=\pi(1-10^{-4})$ for example, the deflection angle mimics the ringdown frequency to one part in $10^{10}$!
Furthermore, we consider the effect of incorporating the QNM’s mirror mode with equal energy, reflecting a more realistic scenario where the SMBHB was in a quasi-circular orbit before the merger. This setup also exhibits the two previously discussed characteristics: an initial rapid decay during the ringdown phase, followed by a late-time power-law tail. Notably, the ringdown phase becomes more pronounced as the metric perturbations are now comparable at both $\theta\sim 0$ and $\theta\sim\pi$.
IV Discussion
This study has explored the dynamic lensing of photons during the ringdown phase following the merger of SMBHBs. By calculating photon trajectories around a relaxing Kerr BH using backward ray tracing, we demonstrate how metric perturbations during the ringdown significantly influence photon paths. In some instances, variations in photon deflection angles closely correspond to the GW ringdown behavior, particularly when photons co-propagate with GWs from the BH to the observer, thereby offering insights into the fascinating dynamics of spacetime.
The visualization of ringdown through photon lensing opens a new avenue for multi-messenger astronomy. While pulsar timing arrays Agazie et al. (2023); Antoniadis et al. (2023); Reardon et al. (2023); Xu et al. (2023), astrometry Braginsky et al. (1990); Book and Flanagan (2011); Moore et al. (2017), binary orbits Blas and Jenkins (2022a, b), and fast radio burst timing Lu et al. (2024) may directly detect ringdown GWs from SMBHB mergers, observations from the Event Horizon Telescope Akiyama et al. (2019, 2022) and future upgrades Ayzenberg et al. (2023), including space missions Haworth et al. (2019); Johnson et al. (2024); Akiyama et al. (2024), could provide visualizations of a ‘Waltz’ between GW and photon co-propagation. Each pixel on the observer plane could reveal additional insights by dissecting the metric perturbations in the vicinity of the relaxing BH.
Clearly, even though we focus on ringdown, their quasi-periodic nature indicates that our results can be taken as indicative of what happens during earlier stages in the life of a binary emitting gravitational radiation.
While our use of backward ray tracing sometimes leads to the disappearance or replacement of a new source at a given pixel, predicting the apparent position shift of a source remains challenging. A future research direction will involve studying photon propagation from a given source point towards the observer in a dynamic spacetime. This endeavor will require a novel approach termed forward ray tracing, presently only feasible by leveraging conserved quantities in Kerr spacetime Zhou et al. (2024).
Acknowledgements.
We thank Maarten van de Meent for helpful discussions and for providing us with useful routines for metric reconstruction. We also extend our thanks to Michael Boyle for his insightful discussions on spin-weighted spherical harmonics. Z.Z. acknowledges financial support from China Scholarship Council (No. 202106040037). V.C. is a Villum Investigator and a DNRF Chair. We acknowledge support by VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. We acknowledge financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101007855 and No 101131233. Y.C. acknowledges the support of the Rosenfeld foundation and the European Consortium for Astroparticle Theory in the form of an Exchange Travel Grant.
Supplemental Material: Lensing Tomography of Black Hole Ringdown
Appendix A Metric Perturbations During Ringdown
A.1 Newman-Penrose formalism
For completeness, we review the Newman-Penrose (NP) formalism in this section, adopting the convention from Refs. van de Meent (2016); Spiers et al. (2023); Pound and Wardell (2020). Note the overall sign difference in some NP quantities compared to the common convention, as seen, for example, in Ref. Teukolsky (1973). This ensures that the background values remain consistent with cases where the metric signature is (+,,,(+,-,-,-)( + , - , - , - ). In the NP formalism, four null vectors $e_{(\mu)}^{a}$ are defined as
$e_{(\mu)}^{a}=\left\{e_{(0)}^{a},e_{(1)}^{a},e_{(2)}^{a},e_{(3)}^{a}\right\}\equiv\left\{l^{a},n^{a},m^{a},\bar{m}^{a}\right\}$ (1)
where $l^{a}$ and $n^{a}$ are real null vectors, and $m^{a}$ and $\bar{m}^{a}$ are complex null vectors. They satisfy
$l^{a}n_{a}=-1\quad\text{and}\quad m^{a}\bar{m}_{a}=1\,.$ (2)
The metric can be obtained via
$g_{ab}=-2l_{(a}n_{b)}+2m_{(a}\bar{m}_{b)}\,.$ (3)
The directional covariant derivatives are defined as
$\hat{D}\equiv l^{a}\nabla_{a},\,\hat{\Delta}\equiv n^{a}\nabla_{a},\,\hat{\delta}\equiv m^{a}\nabla_{a},\,\hat{\bar{\delta}}\equiv\bar{m}^{a}\nabla_{a}.$ (4)
The Ricci rotation coefficients are given by
$\gamma_{(a)(b)(c)}\equiv g_{\mu\lambda}e_{(a)}^{\mu}e_{(c)}^{\nu}\nabla_{\nu}e_{(b)}^{\lambda}\,,$ (5)
which can be explicitly written as
$\kappa$ $\equiv-\gamma_{(3)(1)(1)}\,,\,\varpi\equiv-\gamma_{(2)(4)(1)}\,,\,\epsilon\equiv-\frac{1}{2}\left(\gamma_{(2)(1)(1)}+\gamma_{(3)(4)(1)}\right),$ (6)
$\tau$ $\equiv-\gamma_{(3)(1)(2)}\,,\,\nu\equiv-\gamma_{(2)(4)(2)}\,,\,\gamma\equiv-\frac{1}{2}\left(\gamma_{(2)(1)(2)}+\gamma_{(3)(4)(2)}\right),$
$\sigma$ $\equiv-\gamma_{(3)(1)(3)}\,,\,\mu\equiv-\gamma_{(2)(4)(3)}\,,\,\beta\equiv-\frac{1}{2}\left(\gamma_{(2)(1)(3)}+\gamma_{(3)(4)(3)}\right),$
$\rho$ $\equiv-\gamma_{(3)(1)(4)}\,,\,\lambda\equiv-\gamma_{(2)(4)(4)}\,,\,\alpha\equiv-\frac{1}{2}\left(\gamma_{(2)(1)(4)}+\gamma_{(3)(4)(4)}\right)\,.$
The formal adjoint $\mathcal{L}^{\dagger}$ of an operator
$\mathcal{L}=e_{(\mu)}^{a}\nabla_{a}\,$ (7)
can be written as Wald (1978); Dias et al. (2009); Cartas-Fuentevilla (1997)
$\mathcal{L}^{\dagger}=-\mathcal{L}-\nabla_{a}e_{(\mu)}^{a}$ (8)
Therefore, the formal adjoint directional derivative operators are
$\hat{D}^{\dagger}$ $=-(\hat{D}+\varepsilon+\bar{\varepsilon}-\rho-\bar{\rho})\,,\,\hat{\Delta}^{\dagger}=-(\hat{\Delta}-\gamma-\bar{\gamma}+\mu+\bar{\mu}),$ (9)
$\hat{\delta}^{\dagger}$ $=-(\hat{\delta}+\beta-\bar{\alpha}-\tau+\bar{\varpi})\,,\,\hat{\bar{\delta}}^{\dagger}=-(\hat{\bar{\delta}}+\bar{\beta}-\alpha-\bar{\tau}+\varpi)\,.$
In the Kerr background, we use the Kinnersley tetrad as follows:
$e^{\mu}_{\mathrm{(1)}}=l^{\mu}$ $=\frac{1}{\Delta_{\mathrm{BL}}}\left(r^{2}+a^{2},\Delta_{\mathrm{BL}},0,a\right)\,,\quad e^{\mu}_{\mathrm{(2)}}=n^{\mu}=\frac{1}{2\Sigma_{\mathrm{BL}}}\left(r^{2}+a^{2},-\Delta_{\mathrm{BL}},0,a\right),$ (10)
$e^{\mu}_{\mathrm{(3)}}=m^{\mu}$ $=-\frac{\bar{\rho}}{\sqrt{2}}\left(ia\sin\theta,0,1,\frac{i}{\sin\theta}\right)\,,\quad e^{\mu}_{\mathrm{(4)}}=\bar{m}^{\mu}=\frac{\rho}{\sqrt{2}}\left(ia\sin\theta,0,-1,\frac{i}{\sin\theta}\right),$
where $\rho=-1/(r-ia\cos\theta)$ and the overbar denotes complex conjugation. $\Delta_{\mathrm{BL}}$ and $\Sigma_{\mathrm{BL}}$ are defined as
$\Delta_{\mathrm{BL}}\equiv r^{2}-2Mr+a^{2}\,,\quad\Sigma_{\mathrm{BL}}\equiv r^{2}+a^{2}\cos^{2}\theta\,.$ (11)
The Ricci rotation coefficients can be written explicitly as
$\kappa=\lambda=\nu=\sigma=\epsilon=0\,,$ (12)
and
$\rho=-1/(r-ia\cos\theta)\,,\,\beta=-\bar{\rho}\cot\theta/(2\sqrt{2})\,,\,\varpi=ia\rho^{2}\sin\theta/\sqrt{2}\,,$ (13)
$\tau=-ia\rho\bar{\rho}\sin\theta/\sqrt{2}\,,\,\mu=\rho^{2}\bar{\rho}\Delta/2\,,\,\gamma=\mu+\rho\bar{\rho}(r-M)/2\,,\,\alpha=\varpi-\bar{\beta}\,.$
The Wely scalars are defined as
$\Psi_{0}\equiv C_{(1)(3)(1)(3)}=C_{\mu\nu\rho\sigma}l^{\mu}m^{\nu}l^{\rho}m^{\sigma}\,,\,\Psi_{1}\equiv C_{(1)(2)(1)(3)}=C_{\mu\nu\rho\sigma}l^{\mu}n^{\nu}l^{\rho}m^{\sigma}\,,$ (14)
$\Psi_{2}\equiv C_{(1)(3)(4)(2)}=C_{\mu\nu\rho\sigma}l^{\mu}m^{\nu}\bar{m}^{\rho}n^{\sigma}\,,\,\Psi_{3}\equiv C_{(1)(2)(4)(2)}=C_{\mu\nu\rho\sigma}l^{\mu}n^{\nu}\bar{m}^{\rho}n^{\sigma}\,,$
$\Psi_{4}\equiv C_{(2)(4)(2)(4)}=C_{\mu\nu\rho\sigma}n^{\mu}\bar{m}^{\nu}n^{\rho}\bar{m}^{\sigma}\,,$
where $C_{\mu\nu\rho\sigma}$ is the Weyl tensor. In the Kerr background, we have $\Psi_{0}=\Psi_{1}=\Psi_{3}=\Psi_{4}=0$ and
$\Psi_{2}=\rho^{3}M\,.$ (15)
Other useful equations derived from the Bianchi identities are
$\hat{D}\Psi_{2}=3\rho\Psi_{2}\,,\,\hat{\Delta}\Psi_{2}=-3\mu\Psi_{2}\,,\,\hat{\delta}\Psi_{2}=3\tau\Psi_{2}\,,\,\hat{\bar{\delta}}\Psi_{2}=-3\varpi\Psi_{2}\,.$ (16)
A.2 Metric Reconstruction
Cohen, Chrzanowski, and Kegeles (CCK) formulated a method for reconstructing vacuum metric perturbations within a radiation gauge Chrzanowski (1975); Cohen and Kegeles (1974); Kegeles and Cohen (1979). Here, we briefly review the procedure of CCK reconstruction, which is detailed in Ref. Toomani et al. (2022); Spiers et al. (2023); Green et al. (2020); Campanelli and Lousto (1999); Whiting and Price (2005); Stewart (1979). Many of our calculations heavily depend on the Newman-Penrose (NP) formalism, adhering to the conventions outlined in Appendix A.1.
We begin by introducing four operators: $\hat{\mathcal{E}}$, $\hat{\mathcal{O}}$, $\hat{\mathcal{S}}$, and $\hat{\mathcal{T}}$, each with their respective formal adjoints:
  • The linearized Einstein operator $\hat{\mathcal{E}}$, which is defined as
    ^$\begin{split}\hat{\mathcal{E}}_{ab}(h)=\frac{1}{2}\left[-\nabla^{c}\nabla_{c}h_{ab}-\nabla_{a}\nabla_{b}h_{c}{}^{c}+2\nabla^{c}\nabla_{(a}h_{b)c}+g_{ab}\left(\nabla^{c}\nabla_{c}h_{d}{}^{d}-\nabla^{c}\nabla^{d}h_{cd}\right)\right]\,.\end{split}$ (17)
    The operator $\hat{\mathcal{E}}$ is self-adjoint, as indicated by $\hat{\mathcal{E}}=\hat{\mathcal{E}}^{\dagger}$, and it maps a metric perturbation to its corresponding linearized Einstein tensor.
  • The Teukolsky operator $\hat{\mathcal{O}}$, which can be written as Teukolsky (1973)
    $\hat{\mathcal{O}}_{0}=$ $\left(\hat{D}-3\epsilon+\bar{\epsilon}-4\rho-\bar{\rho}\right)\left(\hat{\Delta}-4\gamma+\mu\right)-\left(\hat{\delta}+\bar{\varpi}-\bar{\alpha}-3\beta-4\tau\right)\left(\hat{\bar{\delta}}+\varpi-4\alpha\right)-3\Psi_{2}\,,$ (18)
    $\hat{\mathcal{O}}_{4}=$ $\left(\hat{\Delta}+3\gamma-\bar{\gamma}+4\mu+\bar{\mu}\right)\left(\hat{D}+4\epsilon-\rho\right)-\left(\hat{\overline{\delta}}-\bar{\tau}+\bar{\beta}+3\alpha+4\varpi\right)\left(\hat{\delta}-\tau+4\beta\right)-3\Psi_{2}\,.$
    Then the Teukolsky equation with spin ±$\pm 2$ can be expressed as
    $\hat{\mathcal{O}}_{0}\Psi_{0}^{(1)}=8\pi T_{0}\,,\,\hat{\mathcal{O}}_{4}\Psi_{4}^{(1)}=8\pi T_{4}\,.$ (19)
    Their separable form, Teukolsky master equation, can be written as
    $\hat{\mathcal{O}}_{s}\psi_{s}=8\pi\Sigma T_{s}\,,$ (20)
    where the spin $s$ is designated as either ±$\pm 2$. For $s=-2$,
    $\psi_{-2}=\rho^{-4}\Psi_{4}^{(1)}\,,\hat{\mathcal{O}}_{-2}=2\Sigma\rho^{-4}\hat{\mathcal{O}}_{4}\rho^{4}\,,T_{-2}=2\rho^{-4}T_{4}\,;$ (21)
    conversely, when $s=+2$,
    $\psi_{+2}=\Psi_{0}^{(1)},\,\hat{\mathcal{O}}_{+2}=2\Sigma\hat{\mathcal{O}}_{0}\,,T_{+2}=2T_{0}\,.$ (22)
  • The operator $\hat{\mathcal{S}}$, which acts on a energy-momentum tensor $T_{ab}$ to yield the source $T_{4}$ of the Teukolsky equation, (Eq. 2.15 in Ref. Teukolsky (1973))
    $\hat{\mathcal{S}}^{ab}=$ $\left(\hat{\Delta}+3\gamma-\bar{\gamma}+4\mu+\bar{\mu}\right)\left[(\hat{\overline{\delta}}-2\bar{\tau}+2\alpha)n^{(a}\bar{m}^{b)}-(\hat{\Delta}+2\gamma-2\bar{\gamma}+\bar{\mu})\bar{m}^{a}\bar{m}^{b}\right]$ (23)
    $+\left(\hat{\overline{\delta}}-\bar{\tau}+\bar{\beta}+3\alpha+4\varpi\right)\left[\left(\hat{\Delta}+2\gamma+2\bar{\mu}\right)n^{(a}\bar{m}^{b)}-\left(\hat{\overline{\delta}}-\bar{\tau}+2\bar{\beta}+2\alpha\right)n^{a}n^{b}\right],$
  • The operator $\hat{\mathcal{T}}$, which acts a metric perturbation $h_{ab}$ to the perturbed Weyl scalar $\Psi_{4}^{(1)}$.
    $\hat{\mathcal{T}}^{ab}=$ $-\frac{1}{2}\left\{\left(\hat{\overline{\delta}}-\bar{\tau}+3\alpha+\bar{\beta}\right)\left(\hat{\overline{\delta}}-\bar{\tau}+2\alpha+2\bar{\beta}\right)n^{a}n^{b}\right.$ (24)
    $+\left(\hat{\Delta}+\bar{\mu}+3\gamma-\bar{\gamma}\right)\left(\hat{\Delta}+\bar{\mu}+2\gamma-2\bar{\gamma}\right)\bar{m}^{a}\bar{m}^{b}$
    $-\left[\left(\hat{\Delta}+\bar{\mu}+3\gamma-\bar{\gamma}\right)\right.\left(\hat{\overline{\delta}}-2\bar{\tau}+2\alpha\right)$
    $\left.\left.+\left(\hat{\overline{\delta}}-\bar{\tau}+3\alpha+\bar{\beta}\right)\left(\hat{\Delta}+2\bar{\mu}+2\gamma\right)\right]n^{(a}\bar{m}^{b)}\right\}$
[…]
Figure 1: Relation of operators for $s=-2$.
As pointed out by Wald Wald (1978), one can find the following operator identity
$\hat{\mathcal{O}}_{4}\hat{\mathcal{T}}=\hat{\mathcal{S}}\hat{\mathcal{E}}\,,$ (25)
and its formal adjoint, $\hat{\mathcal{T}}^{{\dagger}}\hat{\mathcal{O}}^{{\dagger}}_{4}=\hat{\mathcal{E}}\hat{\mathcal{S}}^{{\dagger}}$, is also valid. Then, assuming a complex potential $\Phi$, also known as Herz potential, satisfies the adjoint Teukolsky equation $\hat{\mathcal{O}}^{\dagger}_{4}\Phi=0$, we can consequently have
^$\hat{\mathcal{E}}_{ab}(h)=2\operatorname{Re}(\hat{\mathcal{T}}^{\dagger}_{ab}\hat{\mathcal{O}}^{\dagger}_{4}\Phi)=0$ (26)
Therefore, under the assumption, $h_{ab}$ provides a solution to the linearized Einstein equation, and can be expressed as
$h_{ab}=2\operatorname{Re}\left(\hat{\mathcal{S}}_{ab}^{\dagger}\Phi\right)\,.$ (27)
Chrzanowski introduced two radiation gauges: the ingoing radiation gauge (IRG) and the outgoing radiation gauge (ORG) Chrzanowski (1975). In this paper, we choose ORG, which can be written as
$h_{ab}n^{b}=0=g^{ab}h_{ab}\,.$ (28)
Under this gauge condition, the metric satisfies
$h_{nn}=0=h_{ln}=0=h_{nm}=0=h_{n\bar{m}}=0=h_{m\bar{m}}\,.$ (29)
The Hertz potential $\Phi_{\mathrm{ORG}}$ fulfills the adjoint Teukolsky equation $\mathcal{O}^{\dagger}_{4}\Phi_{\mathrm{ORG}}=0$, where
$\mathcal{O}^{\dagger}_{4}=\left(\hat{D}-3\epsilon+\bar{\epsilon}-\bar{\rho}\right)\left(\hat{\Delta}-4\gamma-3\mu\right)-\left(\hat{\delta}-3\beta-\bar{\alpha}+\bar{\varpi}\right)\left(\hat{\bar{\delta}}-4\alpha-3\varpi\right)-3\Psi_{2}\,.$ (30)
One can verify $\mathcal{O}^{\dagger}_{4}\Phi_{\mathrm{ORG}}=0$ is equivalent to the homogeneous Teukolsky equation for a field with $s=+2$, that is,
$\hat{\mathcal{O}}_{4}^{\dagger}\Phi_{\mathrm{ORG}}=\rho^{-4}\mathcal{O}_{0}\left(\rho^{4}\Phi_{\mathrm{ORG}}\right)=0\,,$ (31)
and the metric component can be written as Dias et al. (2009)
$h_{ab}=$$\left\{n_{(a}\bar{m}_{b)}\left[\left(\hat{\bar{\delta}}+\bar{\beta}-3\alpha+\bar{\tau}+\varpi\right)\left(\hat{\Delta}-4\gamma-3\mu\right)\right.\right.$ (32)
$\left.+\left(\hat{\Delta}-3\gamma-\bar{\gamma}+\mu-\bar{\mu}\right)\left(\hat{\bar{\delta}}-4\alpha-3\varpi\right)\right]$
$-n_{a}n_{b}\left(\hat{\bar{\delta}}-\bar{\beta}-3\alpha+\varpi\right)\left(\hat{\bar{\delta}}-4\alpha-3\varpi\right)$
[…]
where “c.c.” stands for complex conjugate part of the whole object. If we define $\mathring{\Phi}_{\mathrm{ORG}}=\rho^{4}\Phi_{\mathrm{ORG}}$, then the non-zero metric component can be written as Lousto and Whiting (2002); van de Meent and Shah (2015)
$h_{ll}\equiv h_{\mu\nu}l^{\mu}l^{\nu}=\hat{\mathcal{H}}_{ll}^{\dagger}\mathring{\Phi}_{\mathrm{ORG}}+\text{c.c.}\,,\,h_{lm}\equiv h_{\mu\nu}l^{\mu}m^{\nu}=\hat{\mathcal{H}}_{lm}^{\dagger}\mathring{\Phi}_{\mathrm{ORG}},$ (33)
$h_{l\bar{m}}\equiv\bar{h}_{lm}\,,\,h_{mm}\equiv h_{\mu\nu}m^{\mu}m^{\nu}=\hat{\mathcal{H}}_{mm}^{\dagger}\mathring{\Phi}_{\mathrm{ORG}},$
where
^$\hat{\mathcal{H}}_{ll}^{\dagger}$ $=-\rho^{-4}\left(\hat{\bar{\delta}}-3\alpha-\bar{\beta}+5\varpi\right)\left(\hat{\bar{\delta}}-4\alpha+\varpi\right)\,,$ (34)
^$\hat{\mathcal{H}}_{lm}^{\dagger}$ $=-\frac{\rho^{-4}}{2}\left\{\left(\hat{\bar{\delta}}-3\alpha+\bar{\beta}+5\varpi+\bar{\tau}\right)\left(\hat{\Delta}+\mu-4\gamma\right)\right.$
$\left.+\left(\hat{\Delta}+5\mu-\bar{\mu}-3\gamma-\bar{\gamma}\right)\left(\hat{\bar{\delta}}-4\alpha+\varpi\right)\right\}\,,$
^$\hat{\mathcal{H}}_{mm}^{\dagger}$ $=-\rho^{-4}\left(\hat{\Delta}+5\mu-3\gamma+\bar{\gamma}\right)\left(\hat{\Delta}+\mu-4\gamma\right)\,.$
While $\mathring{\Phi}_{\mathrm{ORG}}$ does indeed satisfy the vacuum Teukolsky equation with $s=+2$, the field does not correspond with either the Weyl scalar $\Psi^{\mathrm{(1)}}_{0}$ or $\Psi^{\mathrm{(1)}}_{4}$ originating from the reconstructed metric $h_{ab}$. From Fig. 1, it is clear that $\Psi^{\mathrm{(1)}}_{4}$ can be obtained via a fourth-order ordinary differential equation (ODE) along ingoing null rays Ori (2003); van de Meent and Shah (2015); Spiers et al. (2023)
$\Psi_{4}^{(1)}=\hat{\mathcal{T}}^{ab}h_{ab}\,.$ (35)
We can reduce this circularity condition to what is known as a radial inversion relation Spiers et al. (2023); Kofroň (2020); Aksteiner and Bäckdahl (2019); Deadman and Stewart (2010); Pound and Wardell (2020)
$\Psi_{4}^{(1)}=\frac{1}{2}\left(\hat{\Delta}+3\gamma-\bar{\gamma}\right)\left(\hat{\Delta}+2\gamma-2\bar{\gamma}\right)\left(\hat{\Delta}+\gamma-3\bar{\gamma}\right)\left(\hat{\Delta}-4\bar{\gamma}\right)\bar{\Phi}_{\mathrm{ORG}}\,,$ (36)
which can be solved in vacuum algebraically by applying the separation of variable in next section.
A.2.1 QNM to Hertz potential
In Boyer-Lindquist (BL) coordinates, the Teukolsky operator $\hat{\mathcal{O}}_{s}$ with spin $s$ in Eq. (20) can be written as Teukolsky (1973)
$\displaystyle\hat{\mathcal{O}}_{s}\psi_{s}$ $\displaystyle={\left[\frac{\left(r^{2}+a^{2}\right)^{2}}{\Delta_{\mathrm{BL}}}-a^{2}\sin^{2}\theta\right]\frac{\partial^{2}{}\psi_{s}}{\partial t^{2}}}+\frac{4Mar}{\Delta_{\mathrm{BL}}}\frac{\partial^{2}\psi_{s}}{\partial t\partial\phi}+\left[\frac{a^{2}}{\Delta}-\frac{1}{\sin^{2}\theta}\right]\frac{\partial^{2}\psi_{s}}{\partial\phi^{2}}$
$\displaystyle-\Delta_{\mathrm{BL}}^{-s}\frac{\partial}{\partial r}\left(\Delta_{\mathrm{BL}}^{s+1}\frac{\partial\psi_{s}}{\partial r}\right)-\frac{1}{\sin\theta}\frac{\partial}{\partial\theta}\left(\sin\theta\frac{\partial\psi_{s}}{\partial\theta}\right)-2s\left[\frac{a(r-M)}{\Delta_{\mathrm{BL}}}+\frac{i\cos\theta}{\sin^{2}\theta}\right]\frac{\partial\psi_{s}}{\partial\phi}$
$\displaystyle-2s\left[\frac{M\left(r^{2}-a^{2}\right)}{\Delta_{\mathrm{BL}}}-r-ia\cos\theta\right]\frac{\partial\psi_{s}}{\partial t}+\left(s^{2}\cot^{2}\theta-s\right)\psi_{s}$ (37)
where $\psi_{s}$ is a field with spin $s$. In the vacuum case $T_{s}=0$, we can apply the separation of variables by letting
$\psi_{s}(t,r,\theta,\phi)=\frac{1}{\sqrt{2\pi}}\sum_{l,m,n}{}_{s}Z_{\ell mn}\,{}_{s}S_{\ell mn}(\theta)\,{}_{s}R_{\ell mn}(r)e^{-i\omega_{mn}t+im\phi}\,,$ (38)
With $z=\cos\theta$, ${}_{s}S_{\ell mn}(\theta)$ satisfy
$\partial_{z}\left[\left(1-z^{2}\right)\partial_{z}\left[{}_{s}S_{\ell mn}(\theta)\right]\right]+\left[(cz)^{2}-2csz+s+A^{\ell mn}_{s}(c)-\frac{(m+sz)^{2}}{1-x^{2}}\right]{}_{s}S_{\ell mn}(\theta)=0$ (39)
where $c=a\omega$ and ${}_{s}S_{\ell mn}(\theta)e^{im\phi}$ is the spin-weighted spheroidal harmonics. To normalize the spin-weighted spheroidal harmonics, we apply the following normalization condition
[…]
The radial function $R(r)$ is required to fulfill
[…]
where
[…]
Here ${}_{s}A_{\ell mn}$ is the angular separation constant.
In the following, we will revisit the procedure for obtaining the Hertz potential corresponding to specific Weyl scalar $\Psi^{\mathrm{(1)}}_{0}$ or $\Psi^{\mathrm{(1)}}_{4}$, which is discussed in Ref. Ori (2003); van de Meent and Shah (2015). According to the previously mentioned condition, $\hat{\mathcal{O}}_{+2}\mathring{\Phi}_{\mathrm{ORG}}=0$, we can expand $\mathring{\Phi}_{\mathrm{ORG}}$ in modes
$\mathring{\Phi}_{\mathrm{ORG}}=\frac{1}{\sqrt{2\pi}}\sum_{\ell,m,n}C_{\ell mn}\,{}_{+2}R_{\ell mn}(r)\,{}_{+2}S_{\ell mn}(\theta)e^{im\phi-i\omega_{mn}t}\,,$ (43)
where $C_{\ell mn}$ are constant coefficients. In addition, the Hertz potential $\mathring{\Phi}_{\mathrm{ORG}}$ also satisfies Eq. (36), which can be simplified using the following relationship
$\left(\hat{\Delta}-m\bar{\gamma}+n\gamma\right)f=\Delta_{\mathrm{BL}}^{(n-m)/2}\bar{\rho}^{-m}\rho^{n}\hat{\Delta}\left[\Delta_{\mathrm{BL}}^{(m-n)/2}\bar{\rho}^{m}\rho^{-n}f\right]\,.$ (44)
In this equation, $f=f(t,r,\theta,\phi)$ can represent any function. Then Eq. (36) can be written as Lousto and Whiting (2002); van de Meent and Shah (2015); Ori (2003); Pound and Wardell (2020)
$\frac{1}{32}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{0}^{\dagger}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\bar{\mathring{\Phi}}_{\mathrm{ORG}}\right]=\rho^{-4}\Psi_{4}^{\mathrm{(1)}}\,.$ (45)
Here, the operator $\hat{\mathcal{D}}_{0}^{\dagger}$ is defined as:
$\hat{\mathcal{D}}_{0}^{\dagger}=\partial_{r}-\frac{\left(r^{2}+a^{2}\right)\partial_{t}+a\partial_{\phi}}{\Delta_{\mathrm{BL}}}\,.$ (46)
Note that Eq. (45) can be solved easily on the mode basis given by Eq. (43), using the operator
$\hat{\mathcal{D}}_{mn}^{\dagger}=\partial_{r}+i\frac{K_{mn}}{\Delta_{\mathrm{BL}}}=\partial_{r}+i\frac{\omega_{mn}\left(r^{2}+a^{2}\right)-ma}{\Delta_{\mathrm{BL}}}\,,$ (47)
which act on radial modes. The coefficient can be fixed by considering the asymptotic limit approaching infinity and the horizon. This analysis was previously carried out in Ref. Ori (2003); van de Meent and Shah (2015) for real $\omega$. Due to our choice of gauge (ORG), the results are similar to those in Section III.C of Ref. van de Meent and Shah (2015). Here, we briefly review and extend the procedure to complex $\omega$.
$\frac{1}{32}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{0}^{\dagger}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\bar{\mathring{\Phi}}_{\mathrm{ORG}}\right]=\sum_{\ell,m,n}\frac{\bar{C}_{\ell mn}}{32\sqrt{2\pi}}\Delta_{\mathrm{BL}}^{2}\left(\partial_{r}-i\frac{\bar{K}_{mn}}{\Delta_{\mathrm{BL}}}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\,{}_{+2}\bar{R}_{\ell mn}(r)\,{}_{+2}S_{\ell mn}(\theta)e^{-im\phi+i\bar{\omega}_{mn}t}\right]$ (48)
By relabelling $(m,n)$ to $(-m,-n)$ and employing the identities ${}_{s}S_{\ell mn}(\theta)=(-1)^{s+m}{}_{-s}S_{l-m-n}(\theta)$, ${}_{s}\bar{R}_{\ell mn}={}_{s}R_{l-m-n}$, $\bar{K}_{mn}=-K_{-m-ns}$ and $\bar{\omega}_{mn}=-\omega_{-m-n}$, the equation transforms into
$\frac{1}{32}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{0}^{\dagger}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\bar{\mathring{\Phi}}_{\mathrm{ORG}}\right]=\sum_{\ell,m,n}(-1)^{m}\frac{\bar{C}_{\ell-m-n}}{32\sqrt{2\pi}}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{mn}^{\dagger}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\,{}_{+2}\bar{R}_{\ell mn}(r)\,{}_{-2}S_{\ell mn}(\theta)e^{im\phi-i\omega_{mn}t}\right]$ (49)
When considering Eq. (38), we obtain
${}_{-2}Z_{\ell mn}\,{}_{-2}R_{\ell mn}(r)=\frac{(-1)^{m}}{32}\bar{C}_{\ell-m-n}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{mn}^{\dagger}\right)^{4}\left[\Delta_{\mathrm{BL}}^{2}\,{}_{+2}R_{\ell mn}(r)\right]\,.$ (50)
The asymptotic behavior of ${}_{s}R_{\ell mn}$ is as follows Teukolsky and Press (1974); Cook and Zalutskiy (2014):
$\lim_{r^{*}\rightarrow\infty}{}_{s}R_{\ell mn}(r^{*})\sim\begin{cases}\frac{e^{-i\omega r^{*}}}{r}&:\text{ ingoing wave }\\ \frac{e^{i\omega r^{*}}}{r^{2s+1}}&:\text{ outgoing wave. }\end{cases}$ (51)
$\lim_{r^{*}\rightarrow-\infty}{}_{s}R_{\ell mn}\left(r^{*}\right)\sim\begin{cases}e^{ik_{mn}r^{*}}&:\text{ ingoing (out of }\mathrm{BH})\\ \Delta_{\mathrm{BL}}^{-s}e^{-ik_{mn}r^{*}}&:\text{ outgoing (into BH) }\end{cases}$ (52)
where $k_{mn}=\omega_{mn}-\frac{ma}{2Mr_{+}}$, and $r^{*}$ is the tortoise coordinate which is given by
$r^{*}=r+\frac{2M}{r_{+}-r_{-}}\left[r_{+}\ln\left(\frac{r-r_{+}}{2M}\right)-r_{-}\ln\left(\frac{r-r_{-}}{2M}\right)\right]$ (53)
If $f(r)$ is a solution to the radial Teukolsky equation Eq. (41) with spin $s$, then $\Delta_{\mathrm{BL}}^{s}\bar{f}(r)$ is a solution to the radial Teukolsky equation with spin $-s$. For generic modes with $\omega_{mn}\neq 0$, the asymptotic behavior suggests that this relation swaps physical (retarded) boundary conditions with unphysical (advanced) ones. We denoted the relationship as
${}_{-s}R^{\#}_{\ell mn}\equiv\Delta_{\mathrm{BL}}^{s}\,{}_{s}\bar{R}_{\ell mn}.$ (54)
With this in mind, we have
${}_{-2}Z_{\ell mn}\,{}_{-2}R_{\ell mn}(r)=\frac{(-1)^{m}}{32}\bar{C}_{\ell-m-n}\Delta_{\mathrm{BL}}^{2}\left(\hat{\mathcal{D}}_{mn}^{\dagger}\right)^{4}\left[{}_{-2}R^{\#}_{\ell mn}\right]\,.$ (55)
when we insert the asymptotic behavior into Eq. (55), we can find
$C_{\ell mn}=\frac{2}{\omega_{mn}^{4}}(-1)^{\ell+m}{}_{-2}Z_{\ell mn}$ (56)
Here, we used the following relationship
${}_{-2}\bar{Z}_{lmn}=(-1)^{l}{}_{-2}Z_{l-m-n}\,.$ (57)
Consequently, once we have the solution set $\{\omega,\Psi_{4}^{\mathrm{(1)}}\}$ for the Teukolsky equation with $s=-2$, we are able to reconstruct the first-order metric perturbation $h_{ab}$. After reconstructing the metric, we verify its validity by substituting it into the linearized Einstein equation ^$\hat{\mathcal{E}}_{ab}(h)=0$.
We calculate the quasinormal mode using the qnm package Stein (2019), which employs Leaver’s method to the radial equation and the Cook-Zalutskiy spectral approach to the angular sector Leaver (1986); Cook and Zalutskiy (2014). Furthermore, we use the pseudospectral method to get solutions to the radial Teukolsky equation. Here we briefly illustrate our steps. We convert the asymptotic behavior into coordinate $r$ Cook and Zalutskiy (2014)
$\lim_{r\rightarrow\infty}{}_{s}R_{\ell mn}(r)\sim r^{-1-2s+2i\omega_{mn}M}e^{i\omega_{mn}r}$ (58)
$\lim_{r\rightarrow r_{+}}{}_{s}R_{\ell mn}(r)\sim\left(r-r_{+}\right)^{-s-i\sigma_{+}}$ (59)
where $\sigma_{+}=\frac{2\omega_{mn}Mr_{+}-ma}{r_{+}-r_{-}}$. Then we can define a new radial function ${}_{s}\tilde{R}_{\ell mn}(r)$ by
${}_{s}R_{\ell mn}(r)=\left(r-r_{+}\right)^{-s-i\sigma_{+}}r^{-1-s+2i\omega_{mn}M+i\sigma_{+}}e^{i\omega_{mn}r}{}_{s}\tilde{R}_{\ell mn}(r)\,,$ (60)
where ${}_{s}\tilde{R}_{\ell mn}(r)$ is regular at BH horizon and infinity, and we normalize it by applying
$\lim_{r\to\infty}{}_{s}\tilde{R}_{\ell mn}(r)=1$ (61)
We can expand ${}_{s}\tilde{R}_{\ell mn}(r)$ as a series of Chebyshev polynomials and get ${}_{s}S_{\ell mn}(\theta)$ in terms of spin-weighted spherical functions as in Ref. Cook and Zalutskiy (2014). As a result, we can use our numerical results and evaluate them for any point $(t,r,\theta,\phi)$ of spacetime in the subsequent ray tracing section. To ensure the perturbation metric is flawless, we also verified ^$\hat{\mathcal{E}}_{ab}(h)=0$ numerically.
A.3 Radiated energy
In this section, we estimate the radiated energy of a single mode after truncation. The energy flux is given by Berti et al. (2007)
[…]
Then the total energy emitted can be calculated as
[…]
where $t_{0}=r_{*}(r_{o})$. This total energy satisfies the condition $E_{\mathrm{tot}}/M\lessapprox 3\%$ Berti et al. (2007). Assuming that only the $\ell=m=2$ mode dominates the radiated energy, this relationship leads to
$|{}_{s}Z|\lessapprox 0.24\,.$ (64)
It is important to note that this approach differs slightly from truncating the metric using $H(t-r_{*})$. However, it should provide a sufficiently accurate estimation for our purposes.
Appendix B Backward Ray Tracing
B.1 Local oberver basis
The observation of photons at any given point is dependent on the observer. For the case of a pure Kerr black hole, zero angular momentum observers (ZAMOs) are typically employed. In this section, we extend them to dynamically perturbed Kerr spacetime.
Within the Boyer-Lindquist coordinate system, a ZAMO frame can be chosen and expressed in terms of the coordinate basis $\left\{\partial_{t},\partial_{r},\partial_{\theta},\partial_{\phi}\right\}$ Cunha et al. (2016); Frolov and Novikov (1998); Johannsen (2013); Bardeen (1973) as
$\vec{z}_{(t)}^{\mu}=\frac{g_{\phi\phi}\partial_{t}-g_{\phi t}\partial_{\phi}}{\sqrt{g_{\phi\phi}\left(g_{\phi t}^{2}-g_{\phi\phi}g_{tt}\right)}}\,,\,\vec{z}_{(r)}^{\mu}=\frac{\partial_{r}}{\sqrt{g_{rr}}}\,,\,\vec{z}_{(\theta)}^{\mu}=\frac{\partial_{\theta}}{\sqrt{g_{\theta\theta}}}\,,\,\vec{z}_{(\phi)}^{\mu}=\frac{\partial_{\phi}}{\sqrt{g_{\phi\phi}}}\,.$ (65)
The observer basis $\vec{z}_{(\sigma)}^{\nu}$ has a Minkowski normalization
$g^{(0)}_{\mu\nu}\vec{z}_{(\rho)}^{\mu}\vec{z}_{(\sigma)}^{\nu}=\eta_{(\rho)(\sigma)}\,,$ (66)
with $\eta_{(\rho)(\sigma)}=\operatorname{diag}\{-1,1,1,1\}$. Unlike the unperturbed case, Eq. (66) no longer holds when the unperturbed metric $g^{(0)}_{\mu\nu}$ is replaced by the perturbed metric $g_{\mu\nu}$. To address this issue, we employ a Gram-Schmidt orthogonalization process, using the chosen ZAMO tetrad $\vec{v}_{\rho}=\vec{z}_{(\rho)}^{\mu}$ as the initial set of vectors. The Gram-Schmidt process can be expressed as
$\vec{u}_{0}$ $=\vec{v}_{0}\,,$ (67)
$\vec{u}_{1}$ $=\vec{v}_{1}-\operatorname{proj}_{\vec{u}_{0}}\left(\vec{v}_{1}\right)\,,$
$\vec{u}_{2}$ $=\vec{v}_{2}-\operatorname{proj}_{\vec{u}_{0}}\left(\vec{v}_{2}\right)-\operatorname{proj}_{\vec{u}_{1}}\left(\vec{v}_{2}\right)\,,$
$\vec{u}_{3}$ $=\vec{v}_{3}-\operatorname{proj}_{\vec{u}_{0}}\left(\vec{v}_{3}\right)-\operatorname{proj}_{\vec{u}_{1}}\left(\vec{v}_{3}\right)-\operatorname{proj}_{\vec{u}_{2}}\left(\vec{v}_{3}\right)\,,$
where the projection operator is defined as
$\operatorname{proj}_{\vec{u}}\vec{v}=\frac{\langle\vec{u},\vec{v}\rangle}{\langle\vec{u},\vec{u}\rangle}\vec{u}\,,$ (68)
with the inner product evaluated using the metric $g_{\mu\nu}$. The resulting orthonormalized tetrad, denoted as
$\tilde{\vec{z}}_{(\rho)}^{\mu}=\frac{\vec{u}_{\rho}}{\sqrt{\absolutevalue{\langle\vec{u}_{\rho}\,,\vec{u}_{\rho}\rangle}}}$ (69)
satisfies the following condition
$g_{\mu\nu}\tilde{\vec{z}}_{(\rho)}^{\mu}\tilde{\vec{z}}_{(\sigma)}^{\nu}=\eta_{(\rho)(\sigma)}\,.$ (70)
B.2 Impact Parameters
In the image plane of the observer defined above, each photon is assigned Cartesian coordinates $(\tilde{X},\tilde{Y})$, which represent its impact parameters. These coordinates are proportional to the photon’s observation angles $(\tilde{\alpha},\tilde{\beta})$. Here we follow a convention similar to Ref. Cunha et al. (2016), and define them as:
$\tilde{X}\equiv-r\tilde{\beta},\quad\tilde{Y}\equiv r\tilde{\alpha}\,.$ (71)
Considering the geometry of photon detection, we can express the photon’s 4-momentum in the observer’s reference frame as follows (refer to Fig. 1 in Ref. Cunha et al. (2016)):
$p^{(\phi)}=|\vec{P}|\sin\tilde{\beta}\cos\tilde{\alpha}\,,\,p^{(\theta)}=|\vec{P}|\sin\tilde{\alpha}\,,\,p^{(r)}=|\vec{P}|\cos\tilde{\beta}\cos\tilde{\alpha}\,,$ (72)
where $|\vec{P}|=p^{(t)}$. From this, we can derive the photon’s momentum using
$p^{\mu}=\tilde{\vec{z}}_{(\sigma)}^{\mu}p^{(\sigma)}\,.$ (73)
B.3 Kerr–Schild coordinates
To avoid the numerical problem when a photon gets too close to the pole in the BL coordinate system, we perform ray tracing in the Cartesian Kerr–Schild (KS) coordinate system $\{t_{\mathrm{KS}},x,y,z\}$. The coordinate transformations, which can be found in Ref. Christian and Chan (2021); Mädler and Winicour (2019); Font et al. (1999), are also reviewed in the following.
The Kerr metric in Kerr–Schild coordinates $\{t_{\mathrm{KS}},x,y,z\}$ can be expressed as
$g_{\mu\nu}^{(0)}=\eta_{\mu\nu}+fk_{\mu}k_{\nu}\,,$ (74)
$f=\frac{2Mr^{3}}{r^{4}+a^{2}z^{2}}\,,\,k_{\mu}=\left(1,\frac{rx+ay}{r^{2}+a^{2}},\frac{ry-ax}{r^{2}+a^{2}},\frac{z}{r}\right)\,,$
where $r$ is the same as the radial coordinate in the BL coordinates. The radial coordinate $r$ can be defined in terms of the Cartesian KS coordinates as
$\frac{x^{2}+y^{2}}{r^{2}+a^{2}}+\frac{z^{2}}{r^{2}}=1\,.$ (75)
Then the coordinates $\{r,\theta,\phi_{\mathrm{KS}}\}$ can be written explicitly in terms of Cartesian KS as
$r$ $=\left[\frac{R^{2}-a^{2}+\sqrt{\left(R^{2}-a^{2}\right)^{2}+4a^{2}z^{2}}}{2}\right]^{1/2},$ (76)
$\theta$ $=\arccos{\frac{z}{r}}\,,\,\phi_{\mathrm{KS}}=\tan^{-1}(-ay-rx,ax-ry)+\pi\,.$
where
$R=\sqrt{x^{2}+y^{2}+z^{2}}\,.$ (77)
The function $\tan^{-1}(x,y)$, also known as the two-argument arctangent function, calculates the arctangent of $y/x$ while considering the quadrant in which the point $(x,y)$ lies.
The coordinate transformation from BL coordinates can be written as
[…]
$x$ $=\sqrt{r^{2}+a^{2}}\sin\theta\cos\left[\phi_{\mathrm{KS}}+\arctan(a/r)\right]\,,$
$y$ $=\sqrt{r^{2}+a^{2}}\sin\theta\sin\left[\phi_{\mathrm{KS}}+\arctan(a/r)\right]\,,$
$z$ $=r\cos\theta\,,$
where
$\phi_{\mathrm{KS}}=\phi+a\int\frac{dr}{\Delta_{\mathrm{BL}}}\,.$ (79)
Following Ref. Christian and Chan (2021), we fix the integration constant is fixed by letting $t_{\mathrm{KS}}=t$ and $\phi_{\mathrm{KS}}=\phi$ at the initial radius $r_{o}$, so we have
$t_{\mathrm{KS}}$ $=t+\frac{M}{\sqrt{M^{2}-a^{2}}}\left[r_{+}\ln\left(\frac{r-r_{+}}{r_{o}-r_{+}}\right)-r_{-}\ln\left(\frac{r-r_{-}}{r_{o}-r_{-}}\right)\right]\,,$ (80)
$\phi_{\mathrm{KS}}$ $=\phi+\frac{a}{2\sqrt{M^{2}-a^{2}}}\left[\ln\left(\frac{r-r_{+}}{r_{o}-r_{+}}\right)-\ln\left(\frac{r-r_{-}}{r_{o}-r_{-}}\right)\right]\,.$